# A Calculus of Inheritance

### v4

@@ Appendix A Expressive Asymmetry: Proofs
Theorem A.2 (Forward macro-expressibility).
@@
Each of the five rules of $\mathcal{T}$ replaces one $\lambda$-calculus construct with an inheritance-calculus expression that depends only on the construct’s immediate subexpressions. No rule inspects the surrounding context. ∎
Theorem A.3 (Non-expressibility of inheritance-calculus).
There is no macro-expressible translation from inheritance-calculus to the $\lambda$-calculus.
@@
Let $\mathscr{L}_{1}$ be the inheritance-calculus and let $\mathscr{L}_{0}$ be the sublanguage consisting of mixin trees in the image of the translation $\mathcal{T}$ (Section 4), so that $\mathscr{L}_{0}\subset\mathscr{L}_{1}$. Define the operational equivalence $\cong_{0}$ on $\mathscr{L}_{0}$ via Böhm trees: $\mathcal{T}(M)\cong_{0}\mathcal{T}(N)$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$. Define $\cong_{1}$ on $\mathscr{L}_{1}$ as inheritance-contextual equivalence: $S\cong_{1}S^{\prime}$ iff for every inheritance-calculus context $C[\cdot]$, $C[S]{\Downarrow}\Leftrightarrow C[S^{\prime}]{\Downarrow}$.
By the Böhm Tree Correspondence (Theorems 4.2 and 4.3), $\mathscr{L}_{1}$ conservatively extends $\mathscr{L}_{0}$: for closed $\lambda$-terms $M$, $\mathcal{T}(M){\Downarrow}$ if and only if $M$ has a head normal form.
Define Church booleans via the translation $\mathcal{T}$ (Section 4): $\mathrm{true}=\lambda t.\,\lambda f.\,t$ and $\mathrm{false}=\lambda t.\,\lambda f.\,f$. In de Bruijn notation:
$\mathrm{true}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}$
$\mathrm{false}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!0}.\mathrm{argument}\}\}$
@@
Define Church boolean equality $\mathrm{eq}=\lambda a.\,\lambda b.\,(a\;b)\;(b\;\mathrm{false}\;\mathrm{true})$. In ANF:
$\mathrm{eq}=\lambda a.\,\lambda b.\,\mathbf{let}\;x_{1}=b\;\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;x_{2}=x_{1}\;\mathrm{true}\;\mathbf{in}\;\mathbf{let}\;x_{3}=a\;b\;\mathbf{in}\;x_{3}\;x_{2}$
Applying $\mathcal{T}$ (Section 4):
$\mathrm{eq}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},$
$\quad x_{1}\mapsto\{\uparrow^{\!0}.\mathrm{argument},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\quad x_{2}\mapsto\{x_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
@@
$\quad x_{3}\mapsto\{\uparrow^{\!1}.\mathrm{argument},\;\mathrm{argument}\mapsto\uparrow^{\!0}.\mathrm{argument}\},$
$\quad\_\mathrm{tail}\mapsto\{x_{3}.\mathrm{result},\;\mathrm{argument}\mapsto x_{2}.\mathrm{result}\},$
$\quad\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}\}$
Here $\uparrow^{\!1}.\mathrm{argument}$ in $x_{3}$ reaches the outer $\lambda a$ scope (one $\lambda$ up from $\lambda b$), and $\uparrow^{\!0}.\mathrm{argument}$ reaches $\lambda b$ itself.
In inheritance-calculus, applying a function record to an argument is: inherit the function, override its $\mathrm{argument}$ slot, and project $\mathrm{result}$. Define mixin trees $e_{1}$ and $e_{2}$ that apply $\mathrm{eq}$ to its two arguments via this mechanism:
$e_{1}\;=\;\{$ $\mathrm{eqFalse}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\mathrm{applied}\mapsto\{\mathrm{eqFalse}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
$e_{2}\;=\;\{$ $\mathrm{eqTrue}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\mathrm{applied}\mapsto\{\mathrm{eqTrue}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
That is, $e_{1}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{false}\;\mathrm{false}))$ and $e_{2}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{true}\;\mathrm{true}))$, where $\mathrm{ANF}$ converts a $\lambda$-term to A-normal form by naming every intermediate application (Section 4). Since both $\mathrm{ANF}$ and $\mathcal{T}$ are macro-expressible, their composition is too, so $e_{1},e_{2}\in\mathscr{L}_{0}$. Both $e_{1}.\mathrm{result}$ and $e_{2}.\mathrm{result}$ evaluate to Church $\mathrm{true}$. By Full Abstraction (Theorem 4.3), $\mathrm{BT}(e_{1})=\mathrm{BT}(e_{2})$, so $e_{1}\cong_{0}e_{2}$.
[…]
$C={}$ $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{$
$\qquad\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument},\;\mathrm{secondOperand}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}\}\}$
@@
The de Bruijn indices count scope levels from $\mathrm{repr}$: $\uparrow^{\!0}$ is $\mathrm{repr}$ itself, $\uparrow^{\!1}$ is the $\lambda b$ scope, and $\uparrow^{\!2}$ is the $\lambda a$ scope. Thus $\mathrm{firstOperand}$ projects $\lambda a$’s argument and $\mathrm{secondOperand}$ projects $\lambda b$’s argument.
In the composite scope $\{S,\,C\}$, the definition of $\mathrm{eq}$ gains $\mathrm{repr}$ as a new sibling of $x_{1},x_{2},x_{3}$ inside the $\lambda b$ scope. By open recursion, the de Bruijn references $\uparrow^{\!2}.\mathrm{argument}$ and $\uparrow^{\!1}.\mathrm{argument}$ resolve through the same scope chain as the existing let-bindings. Since $e_{1}$ and $e_{2}$ inherit $\mathrm{eq}$, they also inherit the newly added $\mathrm{repr}$: $e_{1}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{false}$ and $e_{2}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{true}$. These produce structurally distinct mixin trees (one converges to an abstraction selecting its first argument, the other to an abstraction selecting its second), so $e_{1}\not\cong_{1}e_{2}$.

### v6

@@ Appendix C Expressive Asymmetry: Proofs
Lemma C.2 (Composition of macro-expressible translations).
If $\mathcal{E}_{1}$ macro-expresses the constructs of $\mathscr{L}_{1}$ in $\mathscr{L}_{0}$ and $\mathcal{E}_{2}$ macro-expresses the constructs of $\mathscr{L}_{2}$ in $\mathscr{L}_{1}$, then the composition $\mathcal{E}_{1}\circ\mathcal{E}_{2}$ macro-expresses the constructs of $\mathscr{L}_{2}$ in $\mathscr{L}_{0}$.
Proof.
Each construct $F$ of $\mathscr{L}_{2}$ has a fixed syntactic abstraction $A_{F}^{(2)}$ over $\mathscr{L}_{1}$. Each $\mathscr{L}_{1}$-construct $G$ appearing in $A_{F}^{(2)}$ has a fixed syntactic abstraction $A_{G}^{(1)}$ over $\mathscr{L}_{0}$. Substituting $A_{G}^{(1)}$ for every $G$ in $A_{F}^{(2)}$ yields a fixed syntactic abstraction over $\mathscr{L}_{0}$ for $F$, satisfying E4. ∎
Definition C.3 (Language universe).
Let $\mathscr{U}$ be the language whose constructors are the union of the constructors of inheritance-calculus, the constructors of the lazy $\lambda$-calculus ($\lambda$-abstraction, application, variable), and $\mathbf{let}$-binding. The semantics of $\mathscr{U}$ is given by translating every $\mathscr{U}$-program to inheritance-calculus in three steps:
  1. (1)
    Convert every maximal $\lambda$-calculus subexpression (i.e. every subtree whose root and all internal nodes use only $\lambda$-calculus constructors) to A-normal form (Flanagan et al., 1993).
  2. (2)
    Apply the translation $\mathcal{T}$ (Section 4) to every ANF $\lambda$-calculus subexpression.
  3. (3)
    Evaluate the resulting inheritance-calculus program by the rules of Section 3.
When a $\mathscr{U}$-expression interleaves $\lambda$-calculus and inheritance-calculus constructors (e.g. a $\mathbf{let}$-binding whose bound value is an inheritance-calculus expression), the translation is applied compositionally from leaves to root: each $\lambda$-calculus construct is replaced by its inheritance-calculus equivalent via $\mathcal{T}$, treating inheritance-calculus subexpressions as opaque values. Since each rule of $\mathcal{T}$ (Section 4) is a fixed syntactic abstraction depending only on the translated subexpressions, this bottom-up application is well-defined and yields a pure inheritance-calculus program for step (3).
A $\mathscr{U}$-program $P$ converges, $\mathrm{eval}_{\mathscr{U}}(P)$, iff its translation converges. The lazy $\lambda$-calculus and inheritance-calculus are conservative restrictions of $\mathscr{U}$: each is obtained by removing the other’s constructors. For a pure inheritance-calculus program, steps (1) and (2) are vacuous, so $\mathrm{eval}_{\mathscr{U}}$ agrees with the inheritance-calculus semantics. For a pure $\lambda$-calculus program, the composition of steps (1)–(3) agrees with lazy evaluation by Adequacy (Theorem 4.2).
The three-step translation is effective (ANF conversion and $\mathcal{T}$ are syntactic transformations) and the recursive evaluation of inheritance-calculus (Section 3) is a semi-decision procedure, so $\mathrm{eval}_{\mathscr{U}}$ is recursively enumerable, as required by Felleisen’s framework (Felleisen, 1991).
Lemma C.4 ($\mathbf{let}$-binding is macro-expressible in the lazy $\lambda$-calculus).
The translation $\mathbf{let}\;x=e_{1}\;\mathbf{in}\;e_{2}\;\mapsto\;(\lambda x.\,e_{2})\;e_{1}$ is a macro-expressible elimination of $\mathbf{let}$ in the lazy $\lambda$-calculus.
Proof.
The translation replaces each $\mathbf{let}$-binding with a $\lambda$-expression that depends only on the two subexpressions $e_{1}$ and $e_{2}$. This is Proposition 3.7 of Felleisen (Felleisen, 1991). ∎
Lemma C.5 (Equal expressiveness of the lazy $\lambda$-calculus and its ANF fragment).
In $\mathscr{U}$, the lazy $\lambda$-calculus and the ANF lazy $\lambda$-calculus (Section 4) are equally macro-expressive.
Proof.
Every ANF term is a $\lambda$-term, so the embedding of the ANF fragment into the full $\lambda$-calculus is the identity, which is trivially macro-expressible. For the reverse direction, the ANF transformation (Flanagan et al., 1993) consists of two steps, each macro-expressible:
  1. (1)
    Application flattening. Replace each nested application $(e_{1}\;e_{2})\;e_{3}$ with $\mathbf{let}\;x=e_{1}\;e_{2}\;\mathbf{in}\;x\;e_{3}$. The syntactic abstraction is $A(\cdot_{1},\cdot_{2},\cdot_{3})=\mathbf{let}\;x=\cdot_{1}\;\cdot_{2}\;\mathbf{in}\;x\;\cdot_{3}$, depending only on the three subexpressions.
  2. (2)
    Value lifting. Replace each free value $V$ used in an application position with $\mathbf{let}\;x=V\;\mathbf{in}\;x$. The syntactic abstraction is $A(\cdot_{1})=\mathbf{let}\;x=\cdot_{1}\;\mathbf{in}\;x$.
By Lemma C.4, each resulting $\mathbf{let}$-binding is itself macro-expressible in the $\lambda$-calculus. By Lemma C.2, composing these macro-expressible translations yields a macro-expressible translation from the full $\lambda$-calculus to its ANF fragment. ∎
Theorem C.6 (Forward macro-expressibility).
@@
We verify condition E4 of Felleisen (Felleisen, 1991, Definition 3.11): for each $\lambda$-calculus construct $F$ of arity $a$, there must exist a fixed $a$-ary syntactic abstraction $A_{F}(\cdot_{1},\ldots,\cdot_{a})$ over inheritance-calculus such that $\mathcal{T}(F(e_{1},\ldots,e_{a}))=A_{F}(\mathcal{T}(e_{1}),\ldots,\mathcal{T}(e_{a}))$. We check each rule of $\mathcal{T}$ (Section 4):
  1. (1)
    $\lambda$-bound variable ($x$ with de Bruijn index $n^{\prime}$). This is a nullary construct (no subexpressions). The syntactic abstraction is the constant $A=\uparrow^{\!n^{\prime}}.\mathrm{argument}$.
  2. (2)
    Let-bound variable ($x$). Nullary. $A=x.\mathrm{result}$.
  3. (3)
    $\lambda$-abstraction ($\lambda x.\,M$). Unary in the body $M$. $A(\cdot_{1})=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\cdot_{1}\}$.
  4. (4)
    Application let-binding ($\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$). Ternary in $V_{1}$, $V_{2}$, and $M$. $A(\cdot_{1},\cdot_{2},\cdot_{3})=\{x\mapsto\{\cdot_{1},\;\mathrm{argument}\mapsto\cdot_{2}\},\;\mathrm{result}\mapsto\cdot_{3}\}$.
  5. (5)
    Value let-binding ($\mathbf{let}\;x=V\;\mathbf{in}\;M$). Binary in $V$ and $M$. $A(\cdot_{1},\cdot_{2})=\{x\mapsto\{\mathrm{result}\mapsto\cdot_{1}\},\;\mathrm{result}\mapsto\cdot_{2}\}$.
  6. (6)
    Tail call ($V_{1}\;V_{2}$). Binary.
    $A(\cdot_{1},\cdot_{2})=\{\mathrm{tailCall}\mapsto\{\cdot_{1},\;\mathrm{argument}\mapsto\cdot_{2}\},\;\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\}$.
In each case the syntactic abstraction $A_{F}$ is a fixed inheritance-calculus context whose holes are filled by the recursive translations of the subexpressions, satisfying E4. ∎
Theorem C.7 (Nonexpressibility of inheritance).
The lazy $\lambda$-calculus cannot macro-express the inheritance constructors of $\mathscr{U}$ (Definition C.3).
@@
We apply Theorem 3.14(i) of Felleisen (Felleisen, 1991) to the pair $L_{0}=\text{lazy $\lambda$-calculus}$, $L_{1}=\mathscr{U}$. By Definition C.3, $\mathscr{U}=L_{0}+\{$mixin-tree constructors}\} is a conservative extension of $L_{0}$ (Definition 3.2 of Felleisen (Felleisen, 1991)). We verify the four conditions:
  1. (i)
    $L_{0}\subseteq L_{1}$: The syntax of $\mathscr{U}$ includes all $L_{0}$-phrases.
  2. (ii)
    No new $L_{0}$-phrases: The mixin-tree constructors do not introduce new phrases in $L_{0}$.
  3. (iii)
    Disjoint constructors: The mixin-tree constructors are disjoint from the $L_{0}$-constructors.
  4. (iv)
    Semantics preserved: For pure $\lambda$-terms, steps (1) and (2) of Definition C.3 are vacuous and step (3) yields lazy evaluation by Adequacy (Theorem 4.2), so $\mathrm{eval}_{\mathscr{U}}$ agrees with $\mathrm{eval}_{L_{0}}$ on $L_{0}$-phrases.
The theorem requires us to exhibit two $L_{0}$-terms $M$ and $N$ with $M\cong_{0}N$ but $M\not\cong_{1}N$, where $\cong_{0}$ is operational equivalence in $L_{0}$ and $\cong_{1}$ is operational equivalence in $\mathscr{U}$.
Operational equivalence in $L_{0}$
By Böhm tree equivalence (Abramsky, 1990) (Theorem 4.3), for closed $\lambda$-terms $M$ and $N$: $M\cong_{0}N$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$.
Separating pair
@@
Define Church booleans and Church boolean equality: $\mathrm{true}=\lambda t.\,\lambda f.\,t$, $\mathrm{false}=\lambda t.\,\lambda f.\,f$, and $\mathrm{eq}=\lambda a.\,\lambda b.\,(a\;b)\;(b\;\mathrm{false}\;\mathrm{true})$. Let
$M=\mathrm{eq}\;\mathrm{false}\;\mathrm{false}\qquad N=\mathrm{eq}\;\mathrm{true}\;\mathrm{true}.$
Both reduce to Church $\mathrm{true}$ in the lazy $\lambda$-calculus, so $\mathrm{BT}(M)=\mathrm{BT}(N)$ and hence $M\cong_{0}N$.
Distinguishing $\mathscr{U}$-context
We now construct a $\mathscr{U}$-context that separates $M$ and $N$. $M$ and $N$ are syntactically closed $\lambda$-terms; by Definition C.3, their semantics in $\mathscr{U}$ is given by translating them to inheritance-calculus via ANF conversion followed by $\mathcal{T}$. Under the extended ANF convention (Section 4), every free value used in an application is first bound to a name:
$\mathrm{ANF}(M)={}$ $\mathbf{let}\;\mathrm{eq}=\lambda a.\,\lambda b.\,\mathbf{let}\;x_{1}=a\;b\;\mathbf{in}\;$
$\qquad\mathbf{let}\;f_{1}=\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;x_{2}=b\;f_{1}\;\mathbf{in}\;$
$\qquad\mathbf{let}\;t_{1}=\mathrm{true}\;\mathbf{in}\;\mathbf{let}\;x_{3}=x_{2}\;t_{1}\;\mathbf{in}\;x_{1}\;x_{3}$
$\mathbf{in}\;\;\mathbf{let}\;f_{1}=\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;\mathrm{eqFalse}=\mathrm{eq}\;f_{1}$
$\quad\;\mathbf{in}\;\mathbf{let}\;f_{2}=\mathrm{false}\;\mathbf{in}\;\mathrm{eqFalse}\;f_{2}$
Applying $\mathcal{T}$:
$\mathcal{T}(\mathrm{ANF}(M))\;=\;\{$ $\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\mathcal{T}(\lambda a.\,\lambda b.\,\cdots)\},$
$f_{1}\mapsto\{\mathrm{result}\mapsto\mathcal{T}(\mathrm{false})\},$
$\mathrm{eqFalse}\mapsto\{\mathrm{eq}.\mathrm{result},\;\mathrm{argument}\mapsto f_{1}.\mathrm{result}\},$
$f_{2}\mapsto\{\mathrm{result}\mapsto\mathcal{T}(\mathrm{false})\},$
$\mathrm{tailCall}\mapsto\{\mathrm{eqFalse}.\mathrm{result},\;\mathrm{argument}\mapsto f_{2}.\mathrm{result}\},$
$\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\;\}$
and symmetrically for $N$ with $\mathrm{true}$ in place of $\mathrm{false}$. Note that $\mathrm{eq}$ is a named child of $\mathcal{T}(\mathrm{ANF}(M))$, bound via a value let-binding. Inside the translation of $\mathrm{eq}$, the inner $\lambda b$ scope contains the let-bindings for $x_{1}$, $f_{1}$, $x_{2}$, $t_{1}$, $x_{3}$, and a tail call. In particular, $\uparrow^{\!1}.\mathrm{argument}$ reaches the outer $\lambda a$ scope and $\uparrow^{\!0}.\mathrm{argument}$ reaches $\lambda b$ itself. [^17] These indices use the condensed convention, counting only $\lambda$-scopes. As noted in the translation footnote (Section 4), each let-binding and tail call introduces an additional scope level in inheritance-calculus; the formal index is correspondingly larger, but the target scope is the same.
The $\mathscr{U}$-context uses inheritance to add a $\mathrm{repr}$ record inside the $\lambda b$ scope of $\mathrm{eq}$ that projects the first operand. Define the context $C[\cdot]$ as:
$C[\cdot]={}$ $\mathbf{let}\;\mathrm{extended}=\{[\cdot],\;\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{$
@@
$\quad\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument}\}\}\}\}\}$
$\mathbf{in}\;\mathrm{extended}.\mathrm{tailCall}.\mathrm{repr}.\mathrm{firstOperand}$
Here $\cdot$] is the hole. Note the two levels of $\mathrm{result}$ nesting in the mixin: the outer $\mathrm{result}$ enters the value let-binding wrapper for $\mathrm{eq}$, and the inner $\mathrm{result}$ enters the $\lambda a$ scope. The second $\mathrm{result}$ then contains the $\lambda b$ scope, so $\mathrm{repr}$ is placed at the $\lambda b$ scope level. The mixin $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\cdots\}\}\}\}$ is an inheritance-calculus expression: it uses only the mixin-tree constructors of $\mathscr{U}$. The outer $\mathbf{let}$-binding and projection are $\lambda$-calculus constructs (Lemma C.4). Thus $C[\cdot]$ is a valid $\mathscr{U}$-context.
@@
The de Bruijn index $\uparrow^{\!2}$ in $\mathrm{repr}$ counts scope levels (in the condensed convention): $\uparrow^{\!0}$ is $\mathrm{repr}$ itself, $\uparrow^{\!1}$ is the $\lambda b$ scope, and $\uparrow^{\!2}$ is the $\lambda a$ scope. So $\mathrm{firstOperand}$ projects $\lambda a$’s argument.
Since $\mathrm{eqFalse}$ (respectively $\mathrm{eqTrue}$) inherits $\mathrm{eq}.\mathrm{result}$, and $\mathrm{tailCall}$ inherits $\mathrm{eqFalse}.\mathrm{result}$ (the $\lambda b$ scope), the $\mathrm{repr}$ record placed by the mixin at the $\lambda b$ scope level propagates through the inheritance chain to $\mathrm{tailCall}$. In $C[M]$: the $\mathrm{eqFalse}$ binding supplies $\mathrm{argument}\mapsto f_{1}.\mathrm{result}$ (i.e. $\mathrm{false}$) to the $\lambda a$ scope, so $\mathrm{firstOperand}$ evaluates to $\mathrm{false}=\lambda t.\,\lambda f.\,f$. In $C[N]$: the $\mathrm{eqTrue}$ binding supplies $\mathrm{argument}\mapsto\mathrm{true}$, so $\mathrm{firstOperand}$ evaluates to $\mathrm{true}=\lambda t.\,\lambda f.\,t$.
Define $D[\cdot]=(\lambda x.\,x\;\Omega\;I)\;C[\cdot]$. Then $\mathrm{eval}_{\mathscr{U}}$ holds for $D[M]$ (since $\mathrm{false}\;\Omega\;I\to I$, which converges) but not for $D[N]$ (since $\mathrm{true}\;\Omega\;I\to\Omega$, which diverges). Therefore $M\not\cong_{1}N$.
Conclusion